OpenStudy (anonymous):

write an equation for {(5, 2), (0, 12), (1, 10), (-2, 16)}. i dont know how to derive the answer y= - 2x + 12.

OpenStudy (anonymous):

e=mc2

OpenStudy (anonymous):

Take slope of any two points, lets say (0,12) and (1,10) (12-10)/(0-1) 2/-1 -2 now we have slope, use point slope form with lets use point (1,10) y-10=-2(x-1) y-10=-2x+2 y=-2x+12

OpenStudy (anonymous):

thank you so much :)